[figure: 9.7 cm]

A NEURAL NETWORK MODEL OF AN ISING SPIN GLASS

K. F. Wilson^{a} and D. J. Goossens^{a} aDepartment of Physics, Monash University Clayton 3168, Australia

The behaviour of an Ising spin glass (S = \frac{1}{2}) with infinite range interactions is modelled using a numerical simulation based on a neural network. Thermodynamic variables are defined on the network, and are found to obey the Thouless-Anderson-Palmer (TAP) theory when the applied field is zero. When a magnetic field is applied along the spin direction, complex field- dependent behaviour results, including a state in which the Edwards-Anderson order parameter is independent of temperature below the critical temperature.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : K. F. Wilson

Address for correspondence : Department of Physics Monash University Clayton 3168, Australia

Email address : kathryn.wilson@sci.monash.edu.au